Hi,
I was having a difficulty with this age word problem. It goes like this:
John is now 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages.
So I came up with the fact that you can place John as 3x, and Sam as just x, but I couldn't get further than that!
Thank You!

Answer by richwmiller(17219)

j=3s
j+5=2(s+5)
3s+5=2s+10
s=5
j=15
check
15+5=2*(5+5)
20=2*10
ok